# Collection: Datasets from AFAR Challenge

### v1

@@ 0.4 How the Dataset can be Reused
The third column of Fig. 5 presents the RSS and the UAV-to-RF source distance as functions of time. A clear distinction can be observed between the DT and real-world measurements. In the DT environment, RSS remains relatively stable, whereas in the real-world data, significant fluctuations occur—even at similar distances—highlighting the complexity and variability of real-world wireless propagation. These variations in real-world RSS are caused by a combination of factors, including probabilistic LOS conditions, multipath reflections, environmental shadowing, deep fades, and signal blockage from the UAV body itself. Notably, while a general trend is visible—RSS tends to increase as the distance to the RF source decreases—the data clearly shows that distance alone does not fully explain the observed RSS behavior. This challenges the assumptions of conventional path loss models, which typically rely on distance as the primary determinant of RSS. The AFAR dataset offers researchers a unique opportunity to analyze and model these real-world effects using time-synchronized measurements of distance, RSS, and environmental context. By comparing DT and physical measurements, researchers can develop more realistic A2G propagation models. This, in turn, supports the design of more robust localization algorithms and helps bridge the gap between simulation and reality in UAV-assisted wireless systems.
@@
Fig. 6 illustrates the UAV’s orientation—roll, pitch, and yaw—over time, providing valuable insights into flight dynamics and body-induced signal shadowing. Since the receive antenna is mounted on the UAV, changes in orientation significantly affect the antenna pattern, leading to variations in received RSS. In [11], these effects were analyzed using the AFAR dataset. The study highlighted that due to the vertical polarization of the receiving antenna, orientation changes cause polarization mismatch and alter antenna gain, impacting signal reception.
The localization error for each team is shown in Fig. 7. Team T-301 achieved the highest accuracy, winning the competition with a final average localization error of 47.8 meters. Team T-288 secured second place with an average error of 66.6 meters, followed by Team T-300 in third place with an error of 68.3 meters. Detailed descriptions of the algorithms used by these top-performing teams—including their localization strategies, trajectory planning approaches, and underlying motivations—are provided in [8, 10]. Implementation details for Team T-328 can be found in [9]. Team T-309 adopted a fixed trajectory designed to sweep the entire search area within the allotted flight time, ensuring systematic coverage and avoiding unobserved regions. They employed a particle filter-based localization algorithm to estimate the RF source location. Localization error at Loc-3 is consistently higher across all teams compared to the other two locations. This is due to the RF source being positioned outside the UAV flight zone, preventing direct flyovers. The AFAR dataset allows researchers to study such challenging scenarios, enabling the development of improved RSS mapping techniques, trajectory designs, and localization algorithms to handle constrained environments better. To address the change in antenna gains due to UAV orientation, the authors in [11] proposed an enhanced two-ray path loss model that incorporates antenna orientation and UAV body shadowing. Using this refined model, they trained a ResNet-based neural network for localization. The enhanced modeling approach significantly reduced the average localization error from 44 meters (achieved by the top-performing team in the AFAR Challenge) to 18.45 meters. This highlights the importance of accounting for orientation-induced antenna effects and encourages the development of more robust algorithms for advancing UAV-assisted wireless communication systems.
@@
The AFAR dataset offers a comprehensive and versatile foundation for advancing research in UAV-assisted wireless communication. It supports a wide range of applications, including A2G channel modeling, realistic path loss analysis, environmental factor characterization, and localization algorithm validation. The dataset enables researchers to quantify and reduce the gap between simulated and physical deployments, accelerating innovation in 5G/B5G connectivity, autonomous aerial systems, and UAV-based localization. Beyond traditional use cases, the dataset’s rich, time-synchronized measurements unlock opportunities in mobility-aware signal prediction, adaptive trajectory planning, and channel coherence analysis.
@@ ACKNOWLEDGEMENTS
The authors would like to thank Unmanned Experts, Galaxy Unmanned Systems, and AnyMile for being student award sponsors. The article authors have declared no conflicts of interest.

### v2

@@ Abstract
The third column of Fig. 5 presents the RSS and the UAV-to-RF source distance as functions of time. A clear distinction can be observed between the DT and real-world measurements. In the DT environment, RSS follows a relatively smooth, distance-dependent trend, whereas the real-world data exhibit significant fluctuations even at similar distances, highlighting the complexity and variability of RF propagation. These variations arise from probabilistic LoS conditions, multipath reflections, environmental shadowing, deep fades, and UAV body-induced blockage. While a general trend of stronger RSS at shorter distances is visible, the real-world data clearly demonstrate that distance alone cannot fully explain signal behavior, challenging the assumptions of conventional path loss models. The DT environment is based on a two-ray propagation model with one LoS and one ground-reflected NLoS path, combined with Gaussian noise. It replays the same UAV trajectories and software as the real testbed, with transmit power, transmitter/receiver gains, and trajectory parameters calibrated to match the physical experiments. However, this simplified model cannot capture additional reflections, hardware imperfections, or dynamic environmental factors observed in practice.
Quantitatively, for Team T-$328$, the overall mean RSS is identical in both domains (1.29–1.29 dB), but the standard deviation (std) is much higher in the real-world data ($18.09$ dB) compared to the DT ($5.53$ dB). Similarly, at a distance of approximately 150150±55 m, the real-world RSS has a mean of 5.92–5.92 dB with a std of $15.17$ dB, while the DT reports $3.16$ dB with a std of only $5.59$ dB. These results highlight that DT captures the large-scale trend, while the real-world data introduces substantial variability due to fading and environmental influences. Therefore, the two domains should be seen as complementary resources: DT data provide scalable, controlled measurements suitable for pretraining and benchmarking, while real-world datasets enable fine-tuning and validation under realistic conditions. Leveraging both jointly is especially valuable for deep learning workflows, such as transfer learning and domain generalization.
@@
Fig. 6 illustrates the UAV’s orientation—roll, pitch, and yaw over time, providing valuable insights into flight dynamics and body-induced signal shadowing. Since the receive antenna is mounted on the UAV, changes in orientation significantly affect the antenna pattern, leading to variations in received RSS. In [11], these effects were analyzed using the AFAR dataset. The study highlighted that due to the vertical polarization of the receiving antenna, orientation changes cause polarization mismatch and alter antenna gain, impacting signal reception. In our setup, both the UAV receiving antenna and the UGV transmitting antenna are vertically polarized.
The localization error for each team is shown in Fig. 7. Team T-301 achieved the highest accuracy, winning the competition with a final average localization error of 47.8 meters. Team T-288 secured second place with an average error of 66.6 meters, followed by Team T-300 in third place with an error of 68.3 meters. The participating teams employed diverse localization strategies. Team T-301, the winner, used a recursive perimeter-sweep algorithm with averaging to mitigate noise, achieving the best overall accuracy [8]. Team T-288 adopted an online Bayesian optimization framework with Gaussian Process (GP) surrogate modeling, leveraging outlier filtering and adaptive waypoint generation [8, 10]. Team T-300 also relied on GP regression with Bayesian optimization, supplemented by a perimeter survey, but its strong simulation results did not generalize as well to the noisier real-world trials [8]. Team T-309 designed a fixed coverage trajectory combined with a particle filter, which proved sensitive to mismatches between assumed path loss models and real-world propagation. Team T-328 applied a least-squares method based on a path loss model using sampled RSS values, offering a simple but less robust approach [9]. Localization error at Loc-3 is consistently higher across all teams compared to the other two locations. This is due to the RF source being positioned outside the UAV flight zone, preventing direct flyovers. These varied strategies demonstrate the complementary trade-offs between robustness, adaptability, and computational complexity, and provide useful insights for future research on UAV-based RF localization.
To address the change in antenna gains due to UAV orientation, the authors in [11] proposed an enhanced two-ray path loss model that incorporates antenna orientation and UAV body shadowing. Using this refined model, they trained a ResNet-based neural network for localization. The enhanced modeling approach significantly reduced the average localization error from 44 meters (achieved by the top-performing team in the AFAR Challenge) to 18.45 meters. This highlights the importance of accounting for orientation-induced antenna effects and encourages the development of more robust algorithms for advancing UAV-assisted wireless communication systems.
@@
The dataset is subject to certain limitations. Environmental factors such as multipath reflections, foliage, and ground clutter can introduce variability in RSS and RSQ. While the data were collected in a rural environment, RSS and RSQ fluctuations may differ in urban deployments where dense clutter, frequent NLoS conditions, and strong reflections are more pronounced.
@@ ACKNOWLEDGEMENTS
The authors thank Unmanned Experts, Galaxy Unmanned Systems, and AnyMile for sponsoring student awards. The article authors have declared no conflicts of interest.
S. Masrur, Ö. Özdemir, and A. Gürses led the manuscript writing and organization. All other authors contributed to editing and review. Five independent teams participated in the AFAR Challenge. Contributors to algorithm development, UAV trajectory planning, and localization methods include S. Masrur, C. Dickerson, G. Reddy, S. Vargas Villar, C.-W. Wong; B. Chatterjee, S. Chaudhari, Z. Li, and Y. Liu; P. Kudyba and H. Sun; J.S. Mandapaka and K. Namuduri; and W. Wang and F. Fund. The AERPAW platform development and field experiment execution were carried out by Ö. Özdemir, A. Gürses, İ. Güvenç, M.L. Sichitiu, R. Dutta, M. Mushi, and T. Zajkowski.